A body travel the first 1/2 of the total distance with velocity v1 and second 1/2 with velocity v2. Calculated the A.V?

Grade:11

2 Answers

Mudit
7 years ago
Let total distance be s. Velocity to travel distance s/2
Piyush Kumar Behera
7 years ago
There is a formula 2/Vavg=1/v1+1/v2
Actually the question is indirectly asking the proof of this.
To calculate assume that the total distance be 2s.
during the first half ,time taken=s/V1 
during the second half ,time taken=s/V2
Vavg=total distance /total time
Vavg=2s/(s/V1s/V2)
=>2/Vavg=1/v1+1/v2
SO if you are really helped by this solution then please approve this solution to encourage me.